Grade 10Trigonometry

If A and B are zeroes of 4x-5x-1. Find the value of A^2*B+A*B^2

Kate
9 Years agoGrade 10

1 Answer

Anusha Misra
9 Years ago
Im assuming the question is A and B are zeroes of 4x^2-5x-1
so we can infer, A+B = -(-5)/4 = 5/4 (sum of roots)
and AB = -1/4 (product of roots)
A^2*B+A*B^2 = AB (A+B) = -1/4* 5/4 = -5/16